### v2

# Hyunjong Ok1,2  Jaeho Lee1 1POSTECH   2HJ AILAB

Abstract
Singing voices contain much richer information than common voices, such as diverse vocal and acoustic characteristics. However, existing open-source audio-text datasets for singing voices capture only a limited set of attributes and lacks acoustic features, leading to limited utility towards downstream tasks, such as style captioning. To fill this gap, we formally consider the task of singing style captioning and introduce S2Cap, a singing voice dataset with comprehensive descriptions of diverse vocal, acoustic and demographic attributes. Based on this dataset, we develop a simple yet effective baseline algorithm for the singing style captioning. The algorithm utilizes two novel technical components: CRESCENDO for mitigating misalignment between pretrained unimodal models, and demixing supervision to regularize the model to focus on the singing voice. Despite its simplicity, the proposed method outperforms state-of-the-art baselines.
[Uncaptioned image]S2Cap: A Benchmark and a Baseline for Singing Style Captioning
Hyunjong Ok1,2  Jaeho Lee1 1POSTECH   2HJ AILAB hyunjong.ok@gmail.com, jaeho.lee@postech.ac.kr
1 Introduction
Following the recent progress in text-to-speech modeling, the task of speaking style captioning has received a great deal of attention (Yamauchi et al., 2024; Ando et al., 2024; Zhu et al., 2024). Here, the goal is to generate a text prompt that describes para-/non-linguistic characteristics of the speaker from the given audio clip, such as pitch, volume, or gender. The extracted information can contribute greatly to advancing the state-of-the-art of style-conditioned speech synthesis by providing a useful basis for evaluating and labeling the speech data (Wang et al., 2018; Guo et al., 2023; Yang et al., 2024b; Leng et al., 2024).
How much information can speaking style captioning models capture from the singing voices? Singing voices contain rich information regarding the musical characteristics, such as timbre, tempo, or musical genre, providing valuable information for the synthesis and conversion of singing voices (Luo et al., 2020; Zhang et al., 2024; Wang et al., 2024). However, existing speaking style captioning models have been trained to extract only non-musical characteristics from the audio and thus may fall suboptimal for such a purpose. In fact, there is even a lack of appropriate benchmarks to evaluate the performance of such models. Although some datasets provide text attributes or prompts paired with the singing voices (Wang et al., 2022; Huang et al., 2021; Wang et al., 2024), they are limited in the scale of the dataset, the number of attributes, or the diversity of singers (see Table 1).
Figure 1: A comparison of the proposed singing style captioning with related captioning tasks. We did not include speaking style captioning results, as there is no open-sourced model available for the task.
To fill this gap, we formally introduce the task of singing style captioning and take the first step toward solving this task. We first present S2Cap (Singing Style Captioning), a singing voice dataset labeled with ten vocal, musical, and demographic attributes; this is much larger than prior work—e.g., Wang et al. (2024) with only three attributes—enabling the trained model to understand detailed and diverse styles of the singing voices.
Table 1: A comparison of the constructed S2Cap dataset with related singing voice datasets.
Dataset # Singers Duration (h) Style Caption # Attributes
NUS-48E Duan et al. (2013) 12 1.9 -
Opencpop Wang et al. (2022) 1 5.3 -
OpenSinger Huang et al. (2021) 66 50.0 -
M4Singer Zhang et al. (2022) 20 29.8 -
Prompt-Singer Wang et al. (2024) 758 306.9 3
S2Cap (Ours) 2,376 262.9 10
Based on the S2Cap dataset, we establish a novel and robust baseline for the singing style captioning. Similar to conventional audio captioning approaches, the proposed framework leverages pretrained audio encoders and text decoders. To align these models, we develop a new strategy, coined CRESCENDO (Connecting Representations for Efficient Sequence-level Captioning with an End-to-end Transformers), which aligns the features of the audio encoder with an auxiliary text encoder. Furthermore, to let the model focus on the vocal track of the audio, our model is trained using a new loss function that uses demixed vocal audio as an additional source of supervision.
Our contributions can be summarized as follows:
  • We propose a new task called singing style captioning, which aims to generate a text prompt capturing the vocal and musical characteristics of the given audio track.
  • We present S2Cap, a comprehensive open-source dataset for singing style captioning with diverse vocal, musical, and demographic attributes.
  • We establish a strong baseline for the singing style captioning, which includes two novel techniques for addressing the representation misalignment and enhancing model performance through vocal demixing supervision.
Figure 2: An illustration of the data generation pipeline of the S2Cap. We start from a base playlist dataset and collect metadata via web scraping. Then, we partition the audio tracks into multiple segments and extract acoustic information from each segment; through these steps, we collect a total of ten attributes (name, age, genre, tag, volume, pitch, gender, timbre, mood, tempo; marked with a subscript) and the singing style description. Finally, we summarize the collected attributes into a single textual prompt using GPT-4o.
2 Related work
Pretrained language and audio models.
Recent advancements in multimodal learning have leveraged pretrained language and audio models to enhance performance across various tasks (Deshmukh et al., 2023; Chu et al., 2024).
Transformer-based architectures have demonstrated remarkable success in NLP, especially when trained on large-scale corpora. While encoder-only models such as BERT (Devlin et al., 2019) excel in representation learning, they are inherently limited in generative tasks, making them unsuitable for captioning. As a result, encoder-decoder architectures (e.g., BART (Lewis et al., 2020)) and autoregressive decoder-based models (e.g., GPT-2 (Radford et al., 2019)) have become the standard choices for text generation tasks, including multimodal captioning (e.g., image captioning, audio captioning).
Pretrained audio models utilize self-supervised learning techniques that have significantly advanced the field of audio representation learning. Models such as Wav2vec 2.0 (Baevski et al., 2020) have been pretrained on large-scale speech datasets to capture rich acoustic features.
For our model and baselines, we construct and integrate audio encoders with text decoders, demonstrating the potential of cross-modal knowledge transfer for improved caption generation.
Captioning tasks and datasets.
The captioning task aims to generate descriptive texts corresponding to an input from diverse modalities.
Various datasets in the visual domain facilitate research for visual captioning. For image captioning, datasets include Flickr30k (Young et al., 2014), MS COCO (Chen et al., 2015), DenseCap (Johnson et al., 2016), and Conceptual Captions (Sharma et al., 2018), each offering diverse annotations for image descriptions. In video captioning, datasets such as MSVD (Chen and Dolan, 2011), MSR-VTT (Xu et al., 2016), and LSMDC (Rohrbach et al., 2017) support research of textual descriptions from sequential visual content.
In the auditory domain, datasets such as AudioCaps (Kim et al., 2019) and Clotho (Drossos et al., 2020) have been used for general audio captioning tasks. Recently, there has been growing interest in speech style captioning, which involves describing characteristics of human speech (e.g., gender, age group, vocal range) using textual descriptions. PromptTTS (Guo et al., 2023) and InstructTTS (Yang et al., 2024b), for instance, construct dedicated datasets to learn rich speaker-style representations and employ Transformer-based architectures to integrate these representations into synthesized speech. PromptVC (Yao et al., 2024) focuses on text-prompted voice conversion, and PromptSpeaker (Zhang et al., 2023) aims to generate speaker embeddings from text prompts. StyleCap (Yamauchi et al., 2024) initiates speech style captioning to get style descriptions from speech.
Building upon this research, we introduce a new task, singing style captioning, which extends captioning beyond generic speech to the domain of singing with a fine-grained dataset S2Cap that contains various attributes and abundantly explains the singing style through text descriptions.
Music source separation.
Music source separation (MSS), also known as demixing, aims to decompose a mixed music audio signal into its constituent sources, such as vocals, drums, and other accompaniment. Recent advances in deep learning have significantly improved MSS performance, leveraging architectures such as CNN, RNN, and transformer-based networks. Band-split RNN (Luo and Yu, 2023) and HT Demucs (Rouard et al., 2023) have emerged as state-of-the-art models. Band-split RNN employs an RNN-based approach that processes different frequency bands separately, enabling effective modeling of harmonic structures and improving the accuracy of MSS. HT Demucs utilizes a transformer-based architecture to enhance separation performance, capturing local and global dependencies within the audio signal.
In our work, we utilize HT Demucs to extract vocal WAV files from music tracks, which serve as input for our data generation and method.
3 [Uncaptioned image]S2Cap
We now introduce S2Cap, a singing style captioning dataset. The S2Cap consists of 12,105 music tracks with 71,215 textual captions that describe singing styles; each track is partitioned into multiple segments, on which we put separate captions. Along with the captions, each segment has annotations on ten attributes (name, age, gender, tags, volume, pitch, gender, timbre, mood, and tempo). The detailed statistics are given in Table 2.
To avoid license issues, we do not directly share the WAV files for the audio tracks. Instead, we provide the code and URLs to download the wav files. More detailed information and statistics about the dataset are in Appendix A.
Table 2: Basic statistics of the S2Cap dataset
Splits # Tracks # Captions # Words/Caption # Total tokens # Duration (h)
Train 8,395 49,325 23.28 1,148,142 181.86
Dev 1,232 7,339 23.23 170,500 27.22
Test 2,478 14,551 23.30 339,089 53.81
Total 12,105 71,215 23.28 1,657,731 262.88
3.1 Construction Pipeline
For the sake of scale, the S2Cap dataset has been constructed by processing an existing musical dataset with an LLM-based pipeline to generate corresponding textual captions (see Fig. 2). As the base dataset, we have used the Melon playlist dataset, a public dataset containing Mel-spectrogram and textual metadata with over 600,000 tracks (Ferraro et al., 2021). Each track in the dataset has been processed as follows.
Metadata collection and data filtering.
First, we collect textual metadata for each track. This is done by scraping from various web sources, e.g., Last.fm tags. Then, we preprocess the audio track and the metadata as follows. (1) As the Mel-spectrogram in the base dataset is of low resolution, we collect higher-quality audio tracks from YouTube. (2) We filter out the audio files with missing or mismatched metadata entries; we have also filtered out the songs from singers born before 1970, as their songs tend to be missing on YouTube.
Acoustic preprocessing.
Before extracting the acoustic features, we extract two versions of the audio track: the original version and the vocal-only version. The vocal-only version is useful for capturing the style of the vocalist separately from the instrumental parts, and the original version comes in handy in capturing the overall mood. The vocal-only version is prepared using a demixing model, namely the HT Demucs (Rouard et al., 2023). For handling the case of multiple singers with different styles, the demixed tracks are further processed with a speaker diarization model [^1] https://huggingface.co/pyannote/speaker-diarization-3.1; the audio clips are partitioned into 30 seconds-long segments for effective processing, following the prior works (Agostinelli et al., 2023; Doh et al., 2023; Kim et al., 2019). Any segment shorter than 5 seconds is discarded. We also apply the same segmentation to the original version of the audio track.
Acoustic information extraction.
Next, we extract acoustic attributes (volume, pitch, gender, timbre, mood, tempo) and style description prompts from each audio segment. This is done with two pretrained models: Qwen-2 Audio (Chu et al., 2024), and a spoken language understanding (SLU) model. [^2] https://github.com/JeremyCCHsu/Python-Wrapper-for-World-Vocoder Qwen-2 is used to generate annotations on four attributes (gender, timbre, mood, tempo) and the singing style description text; gender and timbre are extracted from the vocal-only version of the audio, and others are generated using the original version. SLU is used to generate volume and pitch annotations. Here, both attributes are classified into three categories—“low,” “medium,” and “high”—where the volume is determined based on the root-mean square of the amplitude and the pitch is determined based on the average $F_{0}$.
Figure 3: An overview of the proposed baseline. An end-to-end transformer comprises an AST encoder and a T5 decoder.
Prompt generation.
Finally, the extracted attributes and singing style description are summarized into a single textual prompt (per segment), with GPT-4o333in particular, the GPT-4o-2024-08-06 (Hurst et al., 2024). The input prompt we used is illustrated in Appendix C. The generated prompts are post-processed, using num2words [^4] https://github.com/savoirfairelinux/num2words to convert numbers to words and languagetool [^5] https://github.com/languagetool-org/languagetool to correct grammatical errors.
Data splitting.
We have partitioned the S2Cap dataset into training/development/test sets in the 70%/10%/20% ratio. We split the dataset so that no artist appears in multiple subsets. Additionally, we have balanced the distribution of six acoustic attributes to preserve the statistical consistency across splits.
Omitted details.
See Appendix B for more details, with an illustrative example of how an audio track is processed with our pipeline.
4 Method
We now describe the proposed baseline method for singing style captioning.
Architecture.
In a nutshell, the model consists of a pretrained audio encoder followed by a pretrained text decoder, which is fine-tuned using the proposed S2Cap dataset (see Fig. 3).
More concretely, consider a Mel-spectrogram of an audio track $X\in\mathbb{R}^{T\times F}$ with $T$ temporal frames and $F$ frequency bins. The audio encoder maps this representation into a feature space
$Z=E_{\text{audio}}(X),$ (1)
where $E_{\text{audio}}(\cdot)$ denotes the audio encoder. In particular, we use the AST encoder (Gong et al., 2021). Given this feature $Z$, the text decoder transforms it into a textual sequence $\hat{Y}=[\hat{y}_{1},\hat{y}_{2},\ldots,\hat{y}_{N}]$, which is a sentence that describes the singing style. That is, the text decoder conducts:
$\hat{Y}=D_{\text{text}}(Z),$ (2)
where $D_{\text{text}}(\cdot)$ denotes the text decoder. We use the text decoder that has a paired text encoder $E_{\text{text}}(\cdot)$; in particular, we use T5 (Raffel et al., 2020).
Training.
The model is trained in two steps:
  1. 1.
    Aligning the feature spaces of the audio encoder and text decoder, using the newly proposed method, CRESCENDO (Section 4.1).
  2. 2.
    Jointly fine-tuning the encoder and decoder, with a novel demixing supervision (Section 4.2).
4.1 Aligning with CRESCENDO [Uncaptioned image]
While both pretrained models $E_{\text{audio}},D_{\text{text}}$ are effective when employed separately for each modality, their embeddings may not be aligned well. Indeed, $D_{\text{text}}$ has been trained to process the features generated by a paired $E_{\text{text}}$ (a T5 encoder, in this case), not $E_{\text{audio}}$ (an AST encoder).
To tackle this issue, the proposed CRESCENDO takes a sequence-level similarity learning approach (Grill et al., 2020; Xu et al., 2022). That is, we fine-tune the audio encoder so that their embeddings match the features of the text encoder $E_{\text{text}}$. A technical obstacle here is that the sequence lengths of the audio and textual embeddings are not necessarily identical; we resolve this issue by applying the mean pooling on the audio features (Fig. 4).
More concretely, let $X$ denote the audio drawn from the S2Cap dataset and $Y$ denote the paired textual prompt. We feed each input to the audio and text encoder, respectively, to get the features
$Z_{\text{a}}=E_{\text{audio}}(X),\quad Z_{\text{t}}=E_{\text{text}}(Y).$ (3)
Then, we maximize the cosine similarity between the text features and the mean-pooled audio features. We also apply the stop-grad operation to the $Z_{\text{t}}$, in order to update only the audio encoder. That is, we use the following loss:
$\mathcal{L}_{\text{CRES}}=1-\frac{\text{pool}(Z_{\text{a}})^{\top}\text{sg}(Z_{\text{t}})}{\|\text{pool}(Z_{\text{a}})\|_{2}\cdot\|\text{sg}(Z_{\text{t}})\|_{2}},$ (4)
where $\text{pool}(\cdot),\text{sg}(\cdot),\|\cdot\|_{2}$ denote the mean pooling, stop-gradient, and $\ell_{2}$ norm, respectively.
Figure 4: Align the audio encoder’s latent space with the text encoder through a sequence-level similarity learning technique. A stop-gradient operation is applied to the text encoder.
4.2 Fine-tuning with demixing supervision
After aligning the feature space of the audio encoder with the feature space of the text decoder, we proceed to fine-tuning the whole model on the S2Cap dataset. The training is done with a mixture of two loss functions, which we describe below.
Cross-entropy loss.
First, we use the standard cross-entropy loss for a more accurate generation of textual captions. We adopt the teacher-forcing strategy, i.e., the model receives the true output token from the previous time step as an input for the current time rather than its generated output. Concretely, the loss can be written as:
$\mathcal{L_{\text{CE}}}=-\sum_{n=1}^{N}\log P(y_{n}|y_{1},\ldots,y_{n-1},E_{\text{audio}}(X))$ (5)
where $P(\cdot|\cdot)$ denotes the output probability of the text decoder and $Y=[y_{1},y_{2},\ldots,y_{N}]$ denotes the text prompt paired with audio $X$.
Demixing supervision loss.
The second loss encourages the audio encoder to focus on extracting features from the singing voice itself rather than the sound from the background, such as instrumental accompaniment. For this purpose, we newly developed the demixing supervision, which regularizes the features extracted from the original tracks to be similar to the features from the vocal-only tracks. More concretely, let $X_{\text{voc}}$ be the vocal-only audio from the S2Cap dataset, paired with the original audio $X$. We minimize the Kullback-Leibler (KL) divergence between the features of two audio tracks. That is, we use the loss
$\mathcal{L_{\text{demix}}}=D_{\text{KL}}\left(E_{\text{audio}}(X_{\text{voc}})\|E_{\text{audio}}(X)\right),$ (6)
where $D_{\text{KL}}(\cdot\|\cdot)$ denotes the KL divergence between two embedding distributions.
Final loss. The final loss is the mixture of the cross-entropy and the demixing supervision losses:
$\mathcal{L_{\text{final}}}=\mathcal{L_{\text{CE}}}+\lambda\cdot\mathcal{L_{\text{demix}}},$ (7)
where $\lambda$ is the weight hyperparameter. We have simply used $\lambda=1$ throughout all experiments.
Table 3: Experiment results on the S2Cap test set of our model structure with various audio encoders with T5. For all metrics, the higher, the better.
Encoder Params BLEU4 METEOR ROUGE-$L$ CIDEr SPICE SPIDEr Sentence-BERT
w/ audioset
AST 86M 29.12 31.50 56.25 105.19 43.24 74.21 85.32
w/ music
MERT 95M 21.78 25.30 49.88 58.99 34.16 46.58 81.62
w/ speech
Wav2vec 2.0 base 95M 28.83 31.12 56.37 102.77 44.57 73.67 86.25
HuBERT base 95M 27.64 30.78 55.67 98.07 43.21 70.64 85.52
5 Experiments
Baselines.
We establish comprehensive baselines for our task. We evaluate transformer-based architectures, aligning with our proposed methodology’s framework. We conduct an extensive ablation study across various audio encoder and text decoder combinations. We evaluate five pre-trained models for audio encoding: AST, MERT (LI et al., 2024), Wav2vec 2.0, and HuBERT (Hsu et al., 2021). These are systematically paired with two decoder variants: GPT-2 and BART-base (w/ decoder part only) (Raffel et al., 2020), exploring eight encoder-decoder configurations. Also, we include two specialized models, Prefix-AAC (Kim et al., 2023) and LP-MusicCaps (Doh et al., 2023). They represent the state-of-the-art in related tasks: audio captioning and music captioning.
Table 4: Experiment results on the S2Cap test set of our method and baselines. For all metrics, the higher, the better.
Methods BLEU4 METEOR ROUGE-$L$ CIDEr SPICE SPIDEr Sentence-BERT
w/ GPT2
AST-GPT2 26.56 30.68 52.64 100.73 44.05 72.39 84.18
MERT-GPT2 26.99 31.05 53.11 101.75 45.19 73.47 85.51
Wav2vec 2.0-GPT2 25.99 30.43 52.41