### v3

# S2Cap: A Benchmark and a Baseline for Singing Style Captioning

Abstract.
[…]
Singing style captioning, Dataset pipeline, Audio-to-text model
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Korea.booktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreaisbn: 979-8-4007-2040-6/2025/11doi: 10.1145/3746252.3761614submissionid: 1328ccs: Computing methodologies Artificial intelligence
1. Introduction
Following the recent progress in text-to-speech modeling, the task of speaking style captioning has received a great deal of attention (Yamauchi et al., 2024; Ando et al., 2024; Zhu et al., 2024). Here, the goal is to generate a text prompt that describes para-/non-linguistic characteristics of the speaker from the given audio clip, such as pitch, volume, or gender. The extracted information can contribute greatly to advancing the state-of-the-art of style-conditioned speech synthesis by providing a useful basis for evaluating and labeling the speech data (Wang et al., 2018; Guo et al., 2023; Yang et al., 2024a; Leng et al., 2024).
How much information can speaking style captioning models capture from the singing voices? Singing voices contain rich musical characteristics, such as timbre, tempo, or musical genre, providing valuable information for the synthesis and conversion of singing voices (Luo et al., 2020; Zhang et al., 2024). However, existing speaking style captioning models have been trained to extract only non-musical characteristics from the audio and thus may fall suboptimal for such a purpose. In fact, there is even a lack of appropriate benchmarks to evaluate the performance of such models. Although some datasets provide text attributes or prompts paired with the singing voices (Wang et al., 2022; Huang et al., 2021; Wang et al., 2024), they are limited in the scale of the dataset, the number of attributes, or the diversity of singers (see Table 1).
To fill this gap, we formally introduce the task of singing style captioning and take the first step toward solving this task. We first present S2Cap (Singing Style Captioning), a singing voice dataset labeled with ten vocal, musical, and demographic attributes; this is much larger than prior work—e.g., Wang et al. (2024) with only three attributes—enabling the trained model to understand detailed and diverse styles of the singing voices.
Table 1. A comparison of the constructed S2Cap dataset with related singing voice datasets.
Dataset # Singers Duration (h) Style Caption # Attributes
NUS-48E (Duan et al., 2013) 12 1.9 -
Opencpop (Wang et al., 2022) 1 5.3 -
OpenSinger (Huang et al., 2021) 66 50.0 -
M4Singer (Zhang et al., 2022) 20 29.8 -
Prompt-Singer (Wang et al., 2024) 758 306.9 3
S2Cap (Ours) 2,376 262.9 10
Building upon the S2Cap dataset, we provide comprehensive baselines by evaluating combinations of diverse audio encoders and text decoders alongside state-of-the-art in related tasks: audio captioning and music captioning. Additionally, we propose a novel training objective designed to enhance the model’s focus on vocal information, utilizing demixed vocal audio as auxiliary supervision. This offers a path for future work to achieve better performance.
Figure 1. An illustration of the data generation pipeline of the S2Cap. We start from a base playlist dataset and collect metadata via web scraping. Then, we partition the audio tracks into multiple segments and extract acoustic information from each segment; through these steps, we collect a total of ten attributes (name, age, genre, tag, volume, pitch, gender, timbre, mood, tempo; marked with a subscript) and the singing style description. Finally, we summarize the collected attributes into a single textual prompt using GPT-4o.
2. Related work
Captioning tasks and datasets.
Captioning task aims to generate descriptive texts corresponding to input from diverse modalities.
Various datasets in the visual domain facilitate research for visual captioning. For image captioning, datasets include Flickr30k (Young et al., 2014), MS COCO (Chen et al., 2015), and Conceptual Captions (Sharma et al., 2018), each offering diverse annotations for image descriptions. In video captioning, datasets such as MSVD (Chen and Dolan, 2011), MSR-VTT (Xu et al., 2016), and LSMDC (Rohrbach et al., 2017) support research of textual descriptions from sequential visual content.
In the auditory domain, datasets such as AudioCaps (Kim et al., 2019) and Clotho (Drossos et al., 2020) have been used for general audio captioning tasks. Recently, there has been growing interest in speech style captioning, which involves describing characteristics of human speech (e.g., gender, age group, vocal range) using textual descriptions. PromptTTS (Guo et al., 2023) and InstructTTS (Yang et al., 2024a), for instance, construct dedicated datasets to learn rich speaker-style representations and employ Transformer-based architectures to integrate these representations into synthesized speech. PromptVC (Yao et al., 2024) focuses on text-prompted voice conversion, and PromptSpeaker (Zhang et al., 2023) aims to generate speaker embeddings from text prompts. StyleCap (Yamauchi et al., 2024) initiates speech style captioning to get style descriptions from speech.
Building upon this research, we introduce a new task, singing style captioning, which extends captioning beyond generic speech to the domain of singing with a fine-grained dataset S2Cap that contains various attributes and abundantly explains the singing style through text descriptions.
Music source separation.
Music source separation (MSS), also known as demixing, aims to decompose a mixed music audio signal into its constituent sources, such as vocals, drums, and other accompaniment. Recent advances in deep learning have significantly improved MSS performance, leveraging architectures such as CNN, RNN, and transformer-based networks. Band-split RNN (Luo and Yu, 2023) and HT Demucs (Rouard et al., 2023) have emerged as state-of-the-art models. Band-split RNN employs an RNN-based approach processing different frequency bands separately, enabling effective harmonic modeling and improving MSS accuracy. HT Demucs utilizes transformer architecture to enhance separation performance, capturing local and global dependencies within audio signals.
In our work, we utilize HT Demucs to extract vocal WAV files, which serve as input for our data generation and method.
3. [Uncaptioned image]S2Cap
We now introduce S2Cap, a singing style captioning dataset. The S2Cap consists of 12,105 music tracks with 71,215 textual captions that describe singing styles; each track is partitioned into multiple segments, on which we put separate captions. Along with the captions, each segment has annotations on ten attributes (name, age, genre, tags, volume, pitch, gender, timbre, mood, and tempo). The detailed statistics are given in Table 2.
To avoid license issues, we do not directly share the WAV files for the audio tracks. Instead, we provide the code and URLs to download the wav files.
Table 2. Basic statistics of the S2Cap dataset.
Splits Tracks Captions Words/Caption Total tokens Duration (h)
Train 8,395 49,325 23.28 1,148,142 181.86
Dev 1,232 7,339 23.23 170,500 27.22
Test 2,478 14,551 23.30 339,089 53.81
Total 12,105 71,215 23.28 1,657,731 262.88
3.1. Construction Pipeline
For the sake of scale, the S2Cap dataset has been constructed by processing an existing musical dataset with an LLM-based pipeline to generate corresponding textual captions (see Fig. 1). As the base dataset, we have used the Melon playlist dataset, a public dataset containing Mel-spectrogram and textual metadata with over 600,000 tracks (Ferraro et al., 2021). Although it originates from a Korean music streaming service, “Melon,” it covers a wide range of international music across diverse genres and artists, beyond Korean songs. Each track in the dataset has been processed as follows.
Metadata collection and data filtering.
First, we collect textual metadata for each track by scraping from various web sources, e.g., Last.fm tags. Then, we preprocess the audio track and the metadata as follows. (1) Since the Mel-spectrogram in the base dataset is of low resolution to avoid licensing issues, we collect higher-quality audio tracks from YouTube. (2) We filter out the audio files with missing or mismatched metadata entries; we also filtered out the songs from singers born before 1970, as their songs tend to be missing on YouTube.
Acoustic preprocessing.
Before extracting the acoustic features, we extract two versions of the audio track: the original version and the vocal-only version. The vocal-only version is useful for capturing the style of the vocalist separately from the instrumental parts, and the original version comes in handy in capturing the overall mood. The vocal-only version is prepared using a demixing model, namely the HT Demucs (Rouard et al., 2023). For handling the case of multiple singers with different styles, the demixed tracks are further processed with a speaker diarization model [^1] https://huggingface.co/pyannote/speaker-diarization-3.1; the audio clips are partitioned into 30 seconds-long segments for effective processing, following the prior works (Agostinelli et al., 2023; Doh et al., 2023; Kim et al., 2019). Any segment shorter than 5 seconds is discarded. We also apply the same segmentation to the original version of the audio track.
Acoustic information extraction.
Next, we extract acoustic attributes (volume, pitch, gender, timbre, mood, tempo) and style description prompts from each audio segment. This is done with two pretrained models: Qwen-2 Audio (Chu et al., 2024), and a spoken language understanding (SLU) model. [^2] https://github.com/JeremyCCHsu/Python-Wrapper-for-World-Vocoder Qwen-2 Audio is used to generate annotations on four attributes (gender, timbre, mood, tempo) and the singing style description text; gender and timbre are extracted from the vocal-only version of the audio, and others are generated using the original version. SLU is used to generate volume and pitch annotations. Here, both attributes are classified into three categories—“low,” “medium,” and “high”—where the volume is determined based on the root-mean square of the amplitude and the pitch is determined based on the average $F_{0}$.
Prompt generation.
Finally, the extracted attributes and singing style description are summarized into a single textual prompt (per segment), with GPT-4o333in particular, the GPT-4o-2024-08-06 (Hurst et al., 2024).
Data splitting.
We have partitioned the S2Cap dataset into training/development/test sets in the 70%/10%/20% ratio. We split the dataset so that no artist appears in multiple subsets. Additionally, we have balanced the distribution of six acoustic attributes to preserve the statistical consistency across splits.
Omitted details.
[…]
4. Experiments
Baselines.
We establish comprehensive baselines for our task. We evaluate transformer-based architectures, aligning with our proposed methodology’s framework. We conduct an extensive ablation study across various audio encoder and text decoder combinations. We evaluate four pretrained models for audio encoding: AST, MERT (LI et al., 2024), Wav2vec 2.0, and HuBERT (Hsu et al., 2021). These are systematically paired with two decoder variants: GPT-2 and BART-base (w/ decoder part only) (Raffel et al., 2020), exploring eight encoder-decoder configurations. Also, we include two specialized models, Prefix-AAC (Kim et al., 2023) and LP-MusicCaps (Doh et al., 2023). They represent the state-of-the-art in related tasks: audio captioning and music captioning. All models are finetuned for $20$ epochs with a batch size of $32$, accumulation steps of $2$, weight decay of $2$, and learning rate of $2\times 10^{-5}$. We use beam search with beam size $5$ during inference.
Evaluation.
To evaluate our proposed methods, we employ metrics that are widely used in captioning tasks such as BLEU (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), ROUGE-L (Lin, 2004), CIDEr (Vedantam et al., 2015), SPICE (Anderson et al., 2016), and SPIDEr (Liu et al., 2017). BLEU is a modified n-gram precision metric incorporating a brevity penalty, while ROUGE-L calculates an F-measure based on the longest common subsequence. METEOR enhances the evaluation by considering several factors like stem- and word-level overlap and synonymy. CIDEr employs a geometric mean of n-gram and cosine similarity scores. SPICE focuses on semantic content by parsing scene graphs from captions, and SPIDEr is the average score of SPICE and CIDEr. While the above metrics are valuable for assessing captioning systems, inspired by recent findings (Zhou et al., 2022; Bhosale et al., 2023), they have limits in capturing the semantic meaning in generated captions. So we supplement our evaluation with Sentence-BERT (Reimers and Gurevych, 2019), metrics tailored for improved semantic relatedness, which produces embeddings for calculating sentence-level similarity.
Additional demixing supervision method.
To enhance the model’s ability to effectively represent singing voices against background music, we introduce a novel fine-tuning strategy that incorporates a demixing supervision loss. This approach regularizes the audio encoder to focus on vocal signals. The training is done with a mixture of two loss functions, which we describe below.
Our training objective combines cross-entropy loss for caption generation with our novel demixing supervision loss. The cross-entropy loss employs teacher-forcing strategy, where the model receives the ground-truth token from the previous step as input:
(1) $\mathcal{L_{\text{CE}}}=-\sum_{n=1}^{N}\log P(y_{n}|y_{1},\ldots,y_{n-1},E_{\text{audio}}(X))$
where $P(\cdot|\cdot)$ denotes the output probability of the text decoder and $Y=[y_{1},y_{2},\ldots,y_{N}]$ denotes the text prompt paired with audio $X$. The demixing supervision loss encourages the audio encoder to extract similar features from both original tracks and their vocal-only counterparts by minimizing the KL divergence:
(2) $\mathcal{L_{\text{demix}}}=D_{\text{KL}}\left(E_{\text{audio}}(X_{\text{voc}})\|E_{\text{audio}}(X)\right),$
where $D_{\text{KL}}(\cdot\|\cdot)$ denotes the KL divergence between two embedding.
The final loss is the mixture of the cross-entropy and the demixing supervision:
(3) $\mathcal{L_{\text{final}}}=\mathcal{L_{\text{CE}}}+\lambda\cdot\mathcal{L_{\text{demix}}},$
where $\lambda$ is the weight hyperparameter.
Table 3. Experiment results on the S2Cap test set of our method and baselines. For all metrics, the higher, the better.
Methods BLEU4 METEOR ROUGE-$L$ CIDEr SPICE SPIDEr Sentence-BERT
w/ GPT2
AST-GPT2 26.56 30.68 52.64 100.73 44.05 72.39 84.18
MERT-GPT2 26.99 31.05 53.11 101.75 45.19 73.47 85.51
Wav2vec 2.0-GPT2 25.99 30.43 52.41 94.37 44.22 69.30 85.01
HuBERT-GPT2 25.80 30.56 52.45 95.41 44.62 70.01 85.28
w/ BART
AST-BART 26.47 30.38 52.50 99.42 43.98 71.70 84.31
MERT-BART 26.49 30.43 52.50 99.95 43.92 71.93 84.04
Wav2vec 2.0-BART 25.51 30.01 51.97 93.30 43.81 68.56 84.53
HuBERT-BART 26.18 30.27 52.60 96.45 44.34 70.39 84.76
w/ T5
AST-T5 29.12 31.50 56.25 105.19 43.24 74.21 85.32
MERT-T5 21.78 25.30 49.88 58.99 34.16 46.58 81.62
Wav2vec 2.0-T5 28.83 31.12 56.37 102.77 44.57 73.67 86.25
HuBERT-T5 27.64 30.78 55.67 98.07 43.21 70.64 85.52
SOTA of related works
Prefix-AAC 29.47 31.64 56.84 104.10 44.87 74.48 86.38
+ w/ Demixing supervision 29.70 31.74 56.97 105.70 44.89 75.29 86.45
LP-MusicCaps 28.33 31.06 55.60 102.92 42.61 72.77 84.91
Figure 2. An example from S2Cap. The emotional intensity of the music gradually escalates from the first to the latter half, effectively captured in the generated caption.
5. Results
5.1. Experiment results
Experiments in various baselines.
We conducted experiments using various encoder-decoder combinations and state-of-the-art (SOTA) models from related works, which are shown in Table 3. Additionally, we applied demixing supervision to the best-performing model Prefix-AAC from our baseline experiments, demonstrating improved performance with demixing supervision.
5.2. Data quality assessment
Human evaluation.
To assess the quality of our dataset, we conducted a human evaluation study by comparing captions generated by GPT-4o, Qwen2.5-72B-Instruct (Yang et al., 2024c), and Llama3.3-70B-Instruct (Dubey et al., 2024) models. Three human annotators were given 200 sampled captions with audio and asked to determine which model produced the best outputs. As shown in Fig. 3, GPT-4o consistently outperformed other models, demonstrating superior caption quality.
In addition to comparative evaluation, we further analyzed GPT-4o-generated caption quality by assessing consistency and fluency criteria, scored out of 5 and 3, respectively. Consistency was evaluated based on factual alignment with music audio tracks, while fluency was measured in terms of grammar, spelling, punctuation, word choice, and sentence structure. Three human annotators were given the same 200 sampled captions with audio and asked to evaluate caption quality. As shown in Table 4, GPT-4o achieved an average consistency score of 4.94 and fluency score of 2.97, confirming the high quality of its generated captions.
To check the objectivity, we have conducted a human evaluation of the timbre generated by Qwen-2 Audio, where the subjective terms come from. Specifically, 20 annotators judged whether the generated timbre appropriately matched one of the 20 well-known singers. Timbre accuracy is 0.75, indicating a strong alignment between human perception and Qwen-2 Audio. Recent studies report human-LLM alignment in evaluations typically about 70% (Kolchinski et al., 2019; Yang et al., 2024b), our result demonstrates notably high agreement.
Table 4. The result of human evaluation on the quality of captions generated by S2Cap.
Methods Consistency Fluency Timbre Acc.
Human eval 4.94 2.97 0.75
Figure 3. Human evaluation results comparing generated captions. The numbers in the bar plot indicate the win rate of the model on the left.
Captioning examples.
We show an example of our S2Cap dataset, as illustrated in Fig. 2. These captions effectively capture various attributes of singing, demonstrating high-quality caption generation. In the given example song, the verse part features a soft and emotional vocal, whereas the chorus gradually intensifies, culminating in an explosive emotional peak. Our dataset successfully reflects these dynamic shifts in the same song.
6. Conclusion
We propose a novel task, singing style captioning, which aims to generate textual prompts describing the vocal characteristics of singers from given song inputs. For this task, we developed S2Cap, a comprehensive dataset reflecting diverse vocal attributes, and established a robust baseline method that effectively captures singing voice characteristics. These contributions provide a solid foundation for future research in this emerging field.
Usage of Generative AI
We use for manuscript refinement and code optimization. Additionally, LLMs are utilized into our data generation pipeline.
Acknowledgements.
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00453301), and by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2019-II191906, Artificial Intelligence Graduate School Program(POSTECH)).